Find a counter example to the following universally quantified statements, where the domain for all variables consist of all integers. ∀x, x^2, <=x^3.
1
Expert's answer
2020-11-04T16:34:52-0500

Let us find a counter example to the following universally quantified statement, where the domain for all variables consist of all integers:

"\forall x (x^2\leq x^3)"


Indeed, for the integer number "x=-2" we have that "x^2=4, x^3=-8". Therefore, it is not true that

"x^2\leq x^3".